# “Medium” LMs of Code in the Era of LLMs: Lessons From StackOverflow

### v2

@@ Abstract.
Large pre-trained neural language models have brought immense progress to both natural language processing and software engineering. Models in OpenAI’s GPT series now dwarf Google’s BERT and Meta’s RoBERTa, which previously set new benchmarks on a wide range of NLP applications. These models are trained on massive corpora of heterogeneous data from web crawls, which enables them to learn general language patterns and semantic relationships. However, the largest models are both expensive to train and deploy and are often closed-source, so we lack access to their data and design decisions. We argue that this trend towards large, general-purpose models should be complemented with single-purpose, more modestly sized pre-trained models. In this work, we take StackOverflow (SO) as a domain example in which large volumes of rich aligned code and text data is available. We adopt standard practices for pre-training large language models, including using a very large context size (2,048 tokens), batch size (0.5M tokens) and training set (27B tokens), coupled with a powerful toolkit (Megatron-LM), to train two models: SOBertBase, with 125 million parameters, and SOBertLarge with 762 million parameters, at a budget of just \$374 and \$1600 each. We compare the performance of our models with both the previous SOTA model trained on StackOverflow data exclusively (BERTOverflow, which relied on an approach tailored to natural language processing) as well general-purpose BERT models (BERTBase and BERTLarge) and OpenAI’s ChatGPT on four StackOverflow-specific downstream tasks - question quality prediction, closed question prediction, named entity recognition and obsoletion prediction (a new task we introduce). Not only do our models consistently outperform all baselines, the smaller model is often sufficient for strong results. Both models are released to the public [^1] https://huggingface.co/mmukh/SOBertLarge, [^2] https://huggingface.co/mmukh/SOBertBase. These results demonstrate that pre-training both extensively and properly on in-domain data can yield a powerful and affordable alternative to leveraging closed-source general-purpose models.
Large Language Models,Pre-trained Models, Discriminative models,Stack Overflow
booktitle: arxiv
1. Introduction
Neural Language Models (LMs) have brought about a significant transformation in the domains of Natural Language Processing (NLP) and Software Engineering (SE). In NLP, LMs are utilized for tasks including text classification (Sun et al., 2019), sentiment analysis (Song et al., 2020), and machine translation (Zhu et al., 2020). In SE, LMs treat code as a form of language and are used for tasks such as code completion (Liu et al., 2020), bug localization (Ciborowska and Damevski, 2022), and code summarization (Wang et al., 2020). Language models can be trained from scratch for a given task as long as they are provided with enough data, typically measured in millions of tokens. In recent years, however, it has been found more effective to pre-train models on large amounts of data to learn general patterns from text or code (Gururangan et al., 2020). Following this pre-training, smaller models such as BERT (Devlin et al., 2018) (with fewer than one billion parameters) are typically fine-tuned on downstream tasks while larger ones such as GPT-3 (Brown et al., 2020) (with many billions of parameters) can be prompted with a natural language question and some examples.
Figure 1. Model and dataset sizes of state-of-the-art LLMs.
Following the observation that language model performance scales predictably and steeply with compute budget (Kaplan et al., 2020), recent trends strongly favor larger models: models like GPT-3, with 175B parameters (Brown et al., 2020), and PALM (Chowdhery et al., 2022) with 540B parameters are trained with hundreds of billions of tokens using a fairly simple, text completion training signal. Commonly referred to as Large Language Models (LLMs) , these models gain remarkable capabilities on a host of language tasks when prompted with just a problem description and some examples. Figure 1 shows the relative scale in terms of both parameters and training data of various BERT models, including ours (in bold) compared to GPT-3 (note the log-scale on both axes). [^3] Where dataset size is only provided in tokens, we assume an average token size of 5 bytes. While GPT-3 is likely significantly smaller than state-of-the-art models in the GPT series, OpenAI has not released details on any more recent models. Yet the ease with which LLMs can be used for new tasks comes at a cost: LLMs are typically massive, requiring millions of dollars of compute to train (Hellendoorn and Sawant, 2021), making them nearly impossible to fine-tune on academic budgets. Most LLMs are also closed-source, so we have no access to their learned weights nor insight into their training data. Finally, as we will show in this work, their general-purpose training signal leaves them performing below-par on tasks where sufficient fine-tuning data is available.
The SE community has responded to the trend towards larger language models with code-specific LMs. For example, CodeBERT (Feng et al., 2020) (110M parameters) is based on BERT, a so-called encoder-only model trained to fill in masked-out tokens in an input. CodeT5 (Wang et al., 2021) (comes in 60M, 220M and 770 parameters) is an encoder-decoder model that encode inputs (mainly, entire functions) plus a task description and decodes (essentially, converts) these into outputs, such as labels, repaired functions. PolyCoder (Xu et al., 2022) (2.7B parameters), Codex(Chen et al., 2021) (12B parameters, not open-source) and CodeGen (Nijkamp et al., 2022) (16B parameters) are all decoder-only models trained to generate code in a left-to-right fashion. In this work, we add to this collection of models by training several BERT-style models on StackOverflow (SO). SO has the unique appeal of providing high-quality text (often, explanations) alongside code, making it a key resource for programmers. Generic (non-SO pre-trained) BERT models are frequently and successfully fine-tuned for downstream tasks involving SO data (Bleyl and Buxton, 2022; Biswas et al., 2020; Kou et al., 2023). Several papers even propose to pre-train such models on SO data, including BERTOverflow (Tabassum et al., 2020), seBERT (von der Mosel et al., 2021) and PTM4Tag (He et al., 2022), often finding that such models are even more powerful when fine-tuned for downstream tasks. However, these pre-training efforts tend to be based in NLP conventions, such as treating each sentence in a post as a separate input to the model. As such, they overlook key properties of StackOverflow posts and large language modeling best practices more generally, like the significant benefits of using very large inputs – BertOverflow limits inputs to 512 tokens, just like CodeT5; we use 2,048 in line with GPT-3 and corpus statistics shown in Figure 4. Unlike seBert (von der Mosel et al., 2021), we preserve the code within the text, as they contain information that is valuable to the overall comprehension of the content. Furthermore, we train models with both posts and their associated comments, which often add key information about the answer’s quality over time.
Figure 2. Framework outlining the key steps of our analysis
With this input representation, we train two BERT models on 27 billion tokens of SO data following best-practice guidelines for training large language models: SOBertBase, with 125M parameters, and SOBertLarge, with 762M parameters. While nontrivial in size, these models can be trained in the cloud at a budget of just \$374 and \$1600 respectively.
Figure 3. With this post (ID: 14569223) as an example, this image demonstrates the structure of a question and answer on SO, featuring a question title, question body, answer body, and comments. We also demonstrate which features we use for each of the following four downstream tasks.
@@ 1. Introduction
We fine-tune both SOBert models as well as an array of baselines including two generic BERT models and BertOverflow, as well as ChatGPT queried via few-shot prompting, on four different tasks that measure StackOverflow question & answer understanding: question quality prediction (Annamoradnejad et al., 2022), closed question prediction(Kag, 2013), named entity recognition (Tabassum et al., 2020) and obsoletion prediction. The latter is a new task that we introduce focused on detecting whether answers contain outdated code, a common phenomenon on SO that can cause vulnerability and compatibility issues for those adopting the code without scrutiny (Fischer et al., 2017; Meng et al., 2018; Ragkhitwetsagul et al., 2019). We constructed this dataset by manually annotating about 1,000 SO answers. SOBert consistently, and often widely, outperforms all other compared models on all four tasks, with even the smaller model, SOBertBase frequently yielding state-of-the-art performance. These results highlight the significant promise of pre-training sub-billion parameter language models on in-domain data when such data is rich in detail and abundant. We release our models for public use.
In summary we make the following contributions:
@@
  • We introduce two models, SOBertBase (125M parameters) and SOBertLarge (762M parameters), that were trained on StackOverflow answer and comment text using the Megatron Toolkit. The training process used multiple GPUs and cost a modest budget of \$374 and \$1600 for SOBertBase and SOBertLarge, respectively.
@@
  • We extensively evaluate both SOBert models on four different tasks, which include question quality prediction, closed question prediction, named entity recognition, and obsoletion prediction.
  • Our evaluation involves comparing the performance of our models to two variants of BERT (BERTBase and BERTLarge), BERTOverflow, and ChatGPT (for three tasks). We demonstrate that SOBert outperforms all other models on all four downstream tasks, and sometimes even SOBertBase outperforms the other compared models in terms of performance.
2. Background
@@
To provide context and prevent ambiguity for the features discussed later, we begin by defining some of the terms we will be using in this paper. We then delve into the background work that has been done in this field.
2.1. How StackOverflow Works
StackOverflow (SO) is a social Q&A platform. SO posts consist of several components, each of which serves a distinct purpose. The primary components of a post are the title, question body, answer body and comments. The title is the headline of the post, and it is meant to provide a brief summary of what the post is about. The question body is the main content of the post, where the user explains the problem they are facing or the question they want to ask. The answer body is where other users can provide their responses to the question asked in the post. Comments are additional remarks that users can add to a post. They can be used to ask for clarification, provide feedback, or suggest improvements to the post. Other components of a post include tags, which are keywords that describe the topic of the post, and votes, which allow users to express their opinion on the quality of the post. Figure 3 provides a visual representation of the various components that make up Stack Overflow (SO) posts. It namely illustrates the different elements, such as the title, question body, answer body and comments that users can interact with when creating or viewing a post on the platform. Users who ask questions are encouraged to both post text describing the problem and include other material such as their (minimal) incorrect code, screenshots, URLs etc. Questions should include between one and five ‘tags’, which identify features like the programming language, library used, or domain of the problem and serve to bring the question to the attention of experienced developers in the corresponding domains. Once a question is posted, developers may respond to with answers, one of which the question-asker may eventually mark as the accepted answer. It is not uncommon for multiple answers to be posted, and upvoted, for a single question; each usually differs in some aspect such as the library (version) used, assumptions about the underlying system & context, or degree of optimization. [^4] Indeed, popular questions frequently have at least one “meta-answer” benchmarking and comparing the solutions proposed in the other answers. Users have the ability to downvote answers or questions as well, to help reduce the visibility of poor quality content. It is not necessary for the accepted answer to have the most upvotes. Along with questions and answers, the website also features a commenting system that allows users to provide feedback or additional information on a particular question or answer. Comments on Stack Overflow can be used for a variety of purposes, such as clarifying a question, suggesting an alternative solution, or pointing out potential issues with an existing answer. They can also be used to ask for further information or to provide additional context that may be helpful in finding a solution to a particular problem.
2.2. Pre-trained Language Models
@@
BERT (Bidirectional Encoder Representations from Transformers) is a type of pre-trained masked language model developed by Google introduced in (Devlin et al., 2018) that can be fine-tuned for various natural language processing tasks. It utilizes the Transformer architecture (Vaswani et al., 2017), which employs a self-attention mechanism to learn a contextual representation of text. It is trained using both the masked language modeling and next sentence prediction objectives on BooksCorpus (800M words) and Wikipedia (2,500M words). Of these two tasks, predicting masked words has proven particularly important as it teaches the model to represent each word in the context of its surrounding words, which helps it learn meaningful representations of the entire document as well. This makes BERT-like models versatile and effective in various downstream NLP tasks such as text classification (Sun et al., 2019; González-Carvajal and Garrido-Merchán, 2020) and named entity recognition (Labusch et al., 2019; Liu and Ritter, 2022).
@@
Two well-trained BERT models were made available by Google. BERTBase-uncased (Devlin et al., 2018) is the smaller model, which has 12 layers of transformer blocks with an internal dimension of 768 for a total of 110 million parameters. The ”uncased” tag means that its input text has been converted to lowercase and all punctuation has been removed, which simplifies learning from a heterogeneous body of text. A second model, BERTLarge (Devlin et al., 2018), has 24 1,024-dimensional transformer layers for a total of 340 million parameters. We match our own Base model, SOBertBase, to the former, while adopting a model with a slightly larger hidden dimension (1,536) than BERTLarge but the same number of layers for SOBertLarge, yielding 762M parameters. We pre-train our models with the masked language modeling objective.
2.3. Modeling StackOverflow data
@@
StackOverflow (SO) frequently releases ‘data dumps’ of publicly available content (questions, answers, comments, etc.). This, combined with its natural alignment of informative text and code, has made it a popular target for SE research. Numerous aspects of software engineering in have been studied on SO by researchers. For instance, one study scrutinized the interactions among developers on the platform (Wang et al., 2013), while another analyzed the themes and patterns in discussions on SO (Barua et al., 2014). Working with StackOverflow data is not without challenges. For one, the quality and structure of answers varies widely from post to post, with many answers being incomplete, containing errors, or lacking context. Another issue is the sheer volume of data on the platform, making it difficult to effectively navigate and identify high-quality task-relevant posts for specific study goals. To mitigate this, studies often focus on a narrow slice of the data, such as investigating the specific inquiries made by software engineers about Android testing on Stack Overflow (Villanes et al., 2017).
@@
Post2Vec, as introduced in Xu et al.’s work (Xu et al., 2021), represents one of the pioneering methods for exploring the generation of distributed representations for Stack Overflow posts. Post2Vec achieved this through the utilization of Convolutional Neural Networks (CNNs) as the primary feature extraction mechanism. At the intersection of language models and SO data, there has been a lot of work in the space of fine-tuning BERT for various tasks such as emotion recognition (Bleyl and Buxton, 2022), sentiment analysis (Biswas et al., 2020), automated summarization (Kou et al., 2023) etc. that involve SO data. Several models have been pre-trained on SO data specifically. He et al.(He et al., 2022) build a tag recommendation framework PTM4TAG for Stack Overflow posts with different BERT-based models. The model seBERT (von der Mosel et al., 2021) is a 340 million parameter model trained on 62.8GB of textual data (they remove all code from the posts) from answer, posts and comments from SO along with more data from GitHub and Jira using the BERTLarge architecture. They use a cased-insensitive 30,522 token WordPiece vocabulary to tokenize their data and restrict inputs to a maximum length of 128 tokens – significantly shorter than most SO posts (Figure 4). Tabassum et al.(Tabassum et al., 2020) released a BERT model pre-trained on StackOverflow data. Their model, BERTOverflow, was pre-trained on 152 million sentences from SO using a case-sensitive 64,000 token WordPiece vocabulary. Training on individual sentences was a common practice in training BERT-like models, but recent work has found that language models perform substantially better when considering a document-level context (Kaplan et al., 2020), albeit at a substantial increase in compute cost. Our findings echo that observation. Tabassum et al. focus on the downstream task of Named Entity Recognition (NER, essentially tagging each word with its syntactic role), on which their model out-performs existing models such as BERTLarge as well as GloVe and ELMo models trained on the same SO data.
@@
We selected BertOverflow as the main specialized baseline in our work. We did not include comparisons to the other models due to differences in model architecture and training data between these models and our proposed SOBert. Specifically, seBERT incorporates additional data sources beyond StackOverflow during pre-training, while PTM4Tag only models the Title, Description, and Code from StackOverflow posts. We omitted a comparison with Post2Vec because its performance has been substantially surpassed by Transformer-based approaches, e.g., as shown in PTM4Tag (He et al., 2022) on a tag prediction task. Instead, we include BERT-Base and BERT-Large as strong, general models, in line with our broader goal of demonstrating that domain-specific pre-trained models can outperform general models. And we include BertOverflow as a representative, capable domain-specific model to highlight how adopting LLM training practices allows for training significantly more powerful domain-specific models.
@@
Compared to these efforts, we pre-train SOBert with 19 GB data presented as just 15 million samples where each sample contains an entire post and all its corresponding comments. We also include all code in each answer so that our model is bimodal in nature. Moreover, we use a SentencePiece tokenizer trained with Byte-Pair Encoding (Kudo and Richardson, 2018), which has the benefit over WordPiece of never labeling tokens as “unknown”. Additionally, SOBert is trained with a maximum sequence length of 2048 based on the empirical length distribution of StackOverflow posts (see Figure 4) and a relatively large batch size of 0.5M tokens (following Kaplan et al.(Kaplan et al., 2020)). Our model performs substantially better than both BertOverflow and BERT models trained on general corpora of text on all downstream tasks (see Section 5).
3. Approach

### v3

# Skill over Scale: The Case for Medium, Domain-Specific Models for SE

@@ Abstract
Recent advancements in AI have sparked a trend in constructing large, generalist language models that handle a multitude of tasks, including many code-related ones. While these models are expensive to train and are often closed-source, they have enjoyed broad adoption because they tend to outperform smaller, domain-specific models of code. In this work, we argue that this is not a foregone conclusion. We show that modestly sized domain-specific models can outperform much larger ones on code labeling tasks, provided they are trained to the same standards. Concretely, we focus on StackOverflow (SO), which offers large volumes of aligned code and text data. We align established best-practices for pre-training large language models with properties of StackOverflow as a data source, especially using a large context window (2,048 tokens), coupled with a powerful toolkit (Megatron-LM) to train two models: SOBertBase[1], with 125M parameters, and SOBertLarge[2] with 762M parameters, at a budget of just \$374 and \$1600 each. We compare the performance of our models with a prior domain-specific model which did not adopt many of these practices (BERTOverflow), as well two general-purpose BERT models (BERTBase and BERTLarge), and two models in OpenAI’s GPT series (GPT-3.5 and GPT-4). We study four labeling tasks: question quality prediction, closed question prediction, named entity recognition and obsoletion prediction. The final task is a new benchmark we introduce, on which we additionally compare SOBert with a fine-tuned CodeLlama and StackLlama (models with 10x more parameters than SOBertLarge). Our models, including the smaller one, consistently outperform all baselines. In contrast, BertOverflow is outperformed by generalist models in most tasks. These results demonstrate that pre-training both extensively and properly on in-domain data can yield a powerful and affordable alternative to leveraging closed-source general-purpose models. Both models are released to the public with over 500 downloads in the last month alone on Hugging Face.
I Introduction
The software development process is changing rapidly, driven by advances in neural Language Models (LMs). While LMs can be trained from scratch on any given dataset, it has been found especially effective to pre-train models on large amounts of data to learn general patterns from text or code [3], and then fine-tune or instruction-tune models for downstream usage [4, 5]. The largest such models, referred to as Large Language Models (LLMs), consistently improve in performance with larger datasets and more parameters [6]. Generalist models, such as GPT-4 [7], therefore train on datasets spanning many languages, domains, and modalities, including source code. These are expensive to train and often closed-source [8], but broadly popular because of their strong performance across many domains. Figure 1 shows the relative scale of various language models, including ours (in bold) compared to GPT-4 (note the log-scale on both axes). [^1] Where dataset size is only reported in terms of tokens, we assume an average token size of 5 bytes. GPT-4 sizes are based on unpublished, but widely repeated claims.
Using code data during pretraining enables LLMs to model code as just another form of language, with significant success [9]. Consequently, LLMs have found many applications in software engineering (SE), including in code completion [10], bug localization [11], and code summarization [12]. This has created a downward pressure on task- and domain-specific models that were previously popular, such as CodeT5 and CodeBERT [13, 14]. In particular, many code-specific resources, including StackOverflow, contain “just” billions of tokens – far too little to train even modest LLMs. It would seem that generalist models are the way forward.
Figure 1: Model and dataset sizes of state-of-the-art LLMs.
Yet the story is not over for domain-specific models of code. While LLMs thrive on generative tasks, they offer less impressive performance on classification tasks in specialized domains. Smaller, BERT-style models thrive on these in large part because they can be fine-tuned easily. LLMs are almost exclusively tuned to new tasks via prompting. In our experiments, for instance, BertOverflow [15], a BERT model trained specifically on StackOverflow content, outperforms GPT-4 on two out of three tasks where they are compared. Surprisingly, though, two generic (data-generalist but small) BERT models (BBU and BLU) mostly outperform GPT-4 and BertOverflow. That might seem to imply that generalist models are still better, as long as fine-tuning is employed, but we conjecture otherwise.
To properly establish whether generalist models should be preferred over domain-specific ones in SE, we need to ensure that the latter are given all the same methodological advantages as the former. The ML community has spent considerable effort developing robust training approaches for LLMs, including extensive scaling laws [16] that relate available compute to ideal model parameter, training configuration, and dataset sizes. This has resulted in training frameworks such as MegatronLM [17] that enable robust training of LMs with, among others, large batch sizes and context windows. We adopt these insights in our work. Going back to the previous case, BBU uses only slightly more data than BertOverflow, but presents that data to the model in chunks of up to 512 tokens. BertOverflow, meanwhile, is trained with sentence-level inputs. StackOverflow posts contain many important inter-sentence relationships, such as between code blocks and explanatory text, and tend to be much longer than 512 tokens (Figure 4). Toolkits for training LLMs now routinely use 2048 or more tokens for context windows.
We demonstrate in this work that the key to producing state-of-the-art models for (at least some) SE tasks lies in training software-specific models with LLM training practices. We train two new BERT-style models on StackOverflow data: SOBertBase and SOBertLarge. We ground their training approach in key characteristics of the dataset and ML best practices. We first collect 19 GB of SO data, where each sample contains an entire post with its comments. We set the context window to a relatively high threshold of 2048 tokens and train a custom vocabulary to maximize the number of characters per token and thus the number of posts that are presented without truncation. In line with established scaling laws [6], we train two relatively small models on this data for multiple epochs (29B tokens seen in total): SOBertBase has 125M parameters, and SOBertLarge has 762M parameters. We adopt hyper-parameters that ensure that both models are trained to high performance on four GPUs. Training these models would cost just ca. \$374 and \$1600 respectively with cloud resources.
@@ I Introduction
We evaluate our models on four StackOverflow labeling tasks. On each task, we fine-tune both the SOBert models as well as three baseline BERT models, GPT-3.5 and GPT-4 prompted in multiple ways, and (for the final task) LoRA [18] fine-tuned LLaMa models with 7B parameters. We study four different tasks that measure SO question & answer understanding: question quality prediction [19], closed question prediction [20], named entity recognition [15] and obsoletion prediction. The latter is a new task that we introduce on detecting whether answers contain outdated code, a common phenomenon on SO that can cause vulnerability and compatibility issues for those adopting the code without scrutiny [21, 22, 23]. SOBert consistently, and often widely, outperforms all other compared models on all tasks, with even the smaller model, SOBertBase frequently yielding state-of-the-art performance.
These results highlight the significant promise of pre-training sub-billion parameter LMs on in-domain data when such data is rich in detail and abundant. We release our models for public use. In summary we make the following contributions:
  • We advocate for a critical approach towards the adoption of LLMs on code-related tasks. Given their cost and proprietary nature, their performance should be judged against diligently trained open-source alternatives.
@@
  • To demonstrate this, we introduce two models, SOBertBase (125M parameters) and SOBertLarge (762M parameters), that were trained on SO answer and comment text using the Megatron Toolkit. The training process used multiple GPUs and cost a modest budget of \$374 and \$1600 for SOBertBase and SOBertLarge, respectively.
@@
  • We extensively evaluate both SOBert models on four different SO tasks: question quality prediction, closed question prediction, named entity recognition, and obsoletion prediction.
  • Compared to three other variants of BERT, GPT-3.5, GPT-4, CodeLlama, StackLlama, and various fine-tuning and prompt engineering strategies, SOBert outperforms all models on all downstream tasks.
Figure 2: Example StackOverflow page (ID: 14569223) demonstrating the structure of a question and answer on SO. Questions have a title and body and may have comments; answers have a body and, optionally, comments. We demonstrate which features are used for each of the four downstream tasks.
II Background
@@
To provide context and reduce ambiguity for the features discussed later, we begin by defining some of the terms we will be using in this paper. We then delve into the background work that has been done in this field.
Our case study, StackOverflow (hereafter abbreviated as SO), is a Q&A platform. Pages on SO consist of the following main components: question title, question body, and or more answers each with a body and (often) comments. The title headlines and summarizes the post, the question body explains the issue, and the answer body provides a response. Comments offer additional remarks for clarification, feedback, or suggestions. Following the SO data terminology, a “Post” is either a question with its title & comments or an answer with its comments (so not the entire Q&A page). Posts additionally come with metadata, such as tags that describe the topic, and votes that gauge post quality. Figure 2 depicts some of these elements in relation to our downstream tasks.
Figure 3: Framework outlining the key steps of our approach
II-A Pre-trained Language Models
@@
BERT (Bidirectional Encoder Representations from Transformers) is a type of pre-trained masked language model developed by Google introduced in [4] that can be fine-tuned for various natural language processing tasks. It utilizes the Transformer architecture [24], which employs a self-attention mechanism to learn a contextual representation of text. It is trained using both the masked language modeling and next sentence prediction objectives on BooksCorpus (800M words) and Wikipedia (2,500M words). Of these two tasks, predicting masked words has proven particularly important as it teaches the model to represent each word in the context of its surrounding words, which helps it learn meaningful representations of the entire document as well. This makes BERT-like models versatile and effective in various downstream NLP tasks such as text classification [25, 26] and named entity recognition [27, 28].
@@
We evaluate SOBert in comparison to seven pre-trained models. BERTBase-uncased [4] is the smaller model, which has 12 layers of transformer blocks with an internal dimension of 768 for a total of 110 million parameters. The ”uncased” tag means that its input text has been converted to lowercase and all punctuation has been removed, which simplifies learning from a heterogeneous body of text. BERTLarge [4], has 24 1,024-dimensional transformer layers for a total of 340 million parameters. We match our own Base model, SOBertBase, to the former, while adopting a model with a slightly larger hidden dimension (1,536) than BERTLarge but the same number of layers for SOBertLarge, yielding 762M parameters. We pre-train our models with the masked language modeling objective.
We also compare SOBert with LLMs from two model families for the obsoletion detection task. These include GPT-4 and GPT-3.5 as closed-source LLMs. We choose CodeLlama and StackLlama as open-source LLMs.
II-B Modeling StackOverflow data
@@
SO frequently releases ‘data dumps’ of publicly available content (questions, answers, comments, etc.). This, combined with its natural alignment of informative text and code, has made it a popular target for SE research. Numerous aspects of SE have been studied on SO by researchers. For instance, one study scrutinized the interactions among developers on the platform [29], while another analyzed the themes and patterns in discussions on SO [30]. Working with SO data is not without challenges. For one, the quality and structure of answers varies widely from post to post, with many answers being incomplete, containing errors, or lacking context. Another issue is the sheer volume of data on the platform, making it difficult to effectively navigate and identify high-quality task-relevant posts for specific study goals. To mitigate this, studies often focus on a narrow slice of the data, such as investigating the specific inquiries made by software engineers about Android testing on SO [31].
@@
Post2Vec, as introduced in Xu et al.’s work [32], represents one of the pioneering methods for exploring the generation of distributed representations for Stack Overflow posts. Post2Vec achieved this through the utilization of Convolutional Neural Networks (CNNs) as the primary feature extraction mechanism. At the intersection of language models and SO data, there has been a lot of work in the space of fine-tuning BERT for various tasks such as emotion recognition [33], sentiment analysis [34], automated summarization [35] etc. that involve SO data. Several models have been pre-trained on SO data specifically. He et al. [36] build a tag recommendation framework PTM4TAG for Stack Overflow posts with different BERT-based models. The model seBERT [37] is a 340 million parameter model trained on 62.8GB of textual data (they remove all code from the posts) from answer, posts and comments from SO along with more data from GitHub and Jira using the BERTLarge architecture. They use a case-insensitive 30,522 token WordPiece vocabulary to tokenize their data and restrict inputs to a maximum length of 128 tokens – significantly shorter than most SO posts (Figure 4). Tabassum et al.[15] released a BERT model pre-trained on SO data. Their model, BERTOverflow, was pre-trained on 152 million sentences from SO using a case-sensitive 64,000 token WordPiece vocabulary. Training on individual sentences was a common practice in training BERT-like models, but recent work has found that language models perform substantially better when considering a document-level context [16], albeit at a substantial increase in compute cost. Our findings echo that observation. Tabassum et al. focus on the downstream task of Named Entity Recognition (NER, essentially tagging each word with its syntactic role), on which their model out-performs existing models such as BERTLarge as well as GloVe and ELMo models trained on the same SO data.
@@
Selection Rationale for Comparative Models: We selected BertOverflow as the main specialized baseline in our work. We did not include comparisons to the other models due to differences in model architecture and training data between these models and our proposed SOBert. Specifically, seBERT incorporates additional data sources beyond SO during pre-training, while PTM4Tag only models the Title, Description, and Code from SO posts. We omitted a comparison with Post2Vec because its performance has been substantially surpassed by Transformer-based approaches, e.g., as shown in PTM4Tag [36] on a tag prediction task. Instead, we include BERT-Base and BERT-Large as strong, general models, in line with our broader goal of demonstrating that domain-specific pre-trained models can outperform general models. And we include BertOverflow as a representative, capable domain-specific model to highlight how adopting LLM training practices allows for training significantly more powerful domain-specific models.
We selected GPT-3.5 and GPT-4 for comparison due to their strong general language understanding capabilities and widespread adoption as closed-source LLMs. We also considered the open-source CodeLLaMA-Instruct [38] and StackLLaMA [39] models, which are pre-trained on code and natural language data, respectively, to leverage their domain-specific knowledge and capabilities. We used the instruction-tuned variant of CodeLLaMA, while StackLLaMA is trained on pairs of questions and answers from Stack Exchange using Direct Preference Optimization (DPO) [40] with the Transformer Reinforcement Learning (TRL) library [41].
@@
Compared to these efforts, we pre-train SOBert with 19 GB data presented as just 15 million samples where each sample contains an entire post and all its corresponding comments. We also include all code in each answer so that our model is bimodal in nature. Moreover, we use a SentencePiece tokenizer trained with Byte-Pair Encoding [42], which has the benefit over WordPiece of never labeling tokens as “unknown”. Additionally, SOBert is trained with a maximum sequence length of 2048, informed by the empirical length distribution of SO posts (see Figure 4) and a relatively large batch size of 0.5M tokens (following Kaplan et al. [16]). Our model performs substantially better than both BertOverflow and BERT models trained on general corpora of text on all downstream tasks (see Section V).
III Approach